NOTICE: This version of the NSF Unidata web site (archive.unidata.ucar.edu) is no longer being updated.
Current content can be found at unidata.ucar.edu.

To learn about what's going on, see About the Archive Site.

[nws-changes] 20081125: TRMM data outage

NOTE: The nws-changes mailing list is no longer active. The list archives are made available for historical reasons.

*UPDATE: 11/25/08, 1725 UTC, 1225 PM, EST

*ESPC is now receiving TRMM data.
The Pacor-A system was experiencing a processing problem that began approximately 330/0300Z. Processing has been moved to their backup system, near realtime data is now available.
*
Details of the Outage:

*ESPC has not received any TRMM (Tropical Rainfall Monitoring Mission) data
since 0302 UTC. There is a problem with a TRMM server at the Goddard Space Flight
center. A system technician is currently working on the problem.

*Data Affected by the Outage:

*TRMM data

*Date and Time of the Outage:
 *
11/24/2008, 0302 UTC, 10:02 PM, EST

*Length of the Outage:
 *
>From 11/24/2008, 0302 UTC, 10:02 PM, EST to 11/25/08, 1725 UTC, 1225 PM, EST.
14 hours and 23 minutes

*Contact Point for Questions:
ESPC Operations  OCL  (Operations Crew Leader),  301-817-3880 *



  • 2008 messages navigation, sorted by:
    1. Thread
    2. Subject
    3. Author
    4. Date
    5. ↑ Table Of Contents
  • Search the nws-changes archives: